# Tug-of-war between idiom’s figurative and literal meanings in LLMs

### v3

@@ 4.3 Identifying attention heads specific to retrieving figurative meanings
Next, we investigated which attention heads are specialized for retrieving an idiom’s figurative meaning, while suppressing its literal compositional interpretation. We knock out attention heads iteratively at the idiom span tokens and measure $\Delta F(s_{a})$, $\Delta L(s_{a})$ to identify the attention heads most impactful for interpretation.
@@
Figure 4: Heatmaps of the (a) $\Delta F(s_{a})$ (b) $\Delta L(s_{a})$ when ablating individual attention heads at the idiom span. Idiomatic heads: Heads those are crucial for retrieving figurative meaning of idiom; $-\Delta F(s_{a})$ and $+\Delta L(s_{a})$.
If a head contributes to retrieving figurative meanings from idioms, we expect that knocking it out would cause the model to lose confidence in the figurative meaning of $s_{a}$, while boosting its literal interpretation. To operationalize this, we define idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ as heads with a large negative $\Delta F(s_{a})$ and positive $\Delta L(s_{a})$. Figure 4 shows that there is indeed a subset of attention heads that simultaneously exhibits large $-\Delta F(s_{a})$ (Figure 4(a)) and $+\Delta L(s_{a})$ (Figure 4(b)). To identify which of these heads are consistently important across instances of $s_{a}$, we performed a nonparametric bootstrap with $B=1000$ resamples to identify, estimating 95% confidence intervals for each head’s contribution to the interpretation. We then rank all heads by, on the one hand, $-\Delta F(s_{a})$, and, on the other hand, $+\Delta L(s_{a})$, and select the top 20 heads highly ranked on both orders, obtaining our list of 20 idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ (selected attention heads listed in Appendix Table 1). Further analyses show that these heads process literal contexts differently compared both to random heads, and to heads that instead focus on compositional semantic processing $\mathcal{H}_{\mathrm{sem}}$ (Appendix B).
@@ 4.4 Causal role of MLP sublayers and idiomatic heads in figurative reading
To validate that the early MLP sublayers and the selected idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ are critical for the figurative reading, we patch their activations on $s_{a}$ into the same components when processing $s_{l}$. As a control, we patch randomly chosen MLP sublayers and heads. We quantify the effect on interpretation by $\Delta F(s_{l}\hookleftarrow s_{a})$ and $\Delta L(s_{l}\hookleftarrow s_{a})$.
ResultsPatching these components boosts the figurative interpretation, as quantified by $\Delta F(s_{l}\hookleftarrow s_{a})$ ($M=0.51,\,SD=0.59$) while suppressing the literal interpretation, $\Delta L(s_{l}\hookleftarrow s_{a})$ ($M=-0.92,\,SD=0.59$). In contrast, patching random components barely affects the figurative interpretation ($\Delta F(s_{l}\hookleftarrow s_{a})$: $M=0.07,\,SD=0.75$) even while suppressing the literal interpretation ($\Delta L(s_{l}\hookleftarrow s_{a})$: $M=-1.83,\,SD=1.39$). Effects on the experiment and the control are significantly different ($p<0.05$). These contrasting effects suggest that literal interpretation depends on scattered non-specialized heads, so perturbing random heads disrupts $\Delta L(s_{l})$. Conversely, idiomatic heads are specialized to enhancing the figurative meaning while suppressing literal interpretations, so that replacing their activations injects a figurative interpretation and amplifies $\Delta F(s_{l})$, while dropping $\Delta L(s_{l})$.
@@ 5 Tracing the interpretation flow
5.1 Locating interpretation representation in token position
After the retrieval process, the model encodes both literal and figurative meanings in its residual stream. We next investigate how these two representations are weighted or integrated in the final prediction. To locate the representation of the meanings, we calculate the mutual nearest-neighbor kernel alignment Huh et al. (2024); Cho et al. (2024) between sentence embeddings of meaning parts from $s_{f}$ and $s_{l}$ (i.e., exclude the ‘X (would)’ and ‘because X was so/a/the’) encoded by BGE M3 Chen et al. (2024) and hidden states from various token positions of idiom sentence $s_{a}$. This metric quantifies the similarity between representations, with higher values indicating greater semantic similarity.
@@ 5.1 Locating interpretation representation in token position
(b) Subsequent token
(c) After subsequent token
@@
ResultsWe plot the kernel alignment at four token positions in Figure 5. We observe that at the subsequent token (Fig. 5(b)), the kernel alignment between the idiom’s hidden states and its figurative meaning surpasses that of the literal meaning after layer 4. This suggests that by this point, the model has begun to favor the figurative interpretation ( 2 selective interpretation step) after retrieving figurative meaning.
@@
Figure 7: Layer-wise interpretation shift after patching in activations from $s_{f},\,s_{l},\,s_{a}^{*}$. The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ is larger than the average difference across all layers; these are statistically significant differences according to a paired $t$-test, $p<0.05$.
To dissociate the computational pathways of processing the figurative and literal interpretations, we employ activation patching at the subsequent token (i.e., ‘because’) across layers, the point where the figurative interpretation is most strongly encoded. We generate three conditions by pairing a source – a figurative context, a literal context, or a different idiom context ($s_{f},\,s_{l},\,s_{a}^{*}$), with a target sentence ($s_{a}$). Then, we patch the layer-wise residual activations at the ‘because’ position from the source condition into the target, see Figure 6 for illustration of expected effect. For each layer $\ell$, we measure the patching causal effect by $\Delta F$ and $\Delta L$. Figure 7 summarizes the shift of interpretation after patching. We additionally patch idiom representations into paraphrased sentences (Appendix 9), which support pathways of interpretations that we find.
ResultsPatching in figurative activations $s_{f}$ into $s_{a}$ removes the intermediate path literal flow but leaves the direct route intact (Figure 6(a)). Figure 7(a) shows that this slightly increases the $\Delta F(s_{a}\hookleftarrow s_{f})$ by up to $0.01$ for patching in mid-layers 4-10, but doesn’t affect the $\Delta L(s_{a}\hookleftarrow s_{f})$, which remains zero. This indicates that for the mid-layers at the ‘because’ token position, the model’s residual stream is the predominant channel for the figurative interpretation ( 3-1 figurative path). That is, we argue that there’s a different path for the literal interpretation ( 3-2 compositional semantic direct path) to convey its information to the final prediction, directly bypassing the intermediate path. This claim is supported by an attention-edge knockout experiment, where we prevent the last token from attending to the idiom tokens, and observe a drop in $\Delta L(s_{a})$ at layer 11 and 13 while $\Delta F(s_{a})$ remains near zero (see Appendix D).
@@ 5.2 Analyzing layer-wise competing interpretation flow through activation patching
We next patch the intermediate path by activations from the literal paraphrase $s_{l}$. This removes the intermediate path figurative flow (Figure 6(b)). Figure 7(b) shows a substantial increase in the $\Delta L(s_{a}\hookleftarrow s_{l})$, by $0.22$ in mid-layers (5-10). At the same time, $\Delta F(s_{a}\hookleftarrow s_{l})$ drops sharply at intermediate layers, which indicates that the figurative interpretation is severely disrupted.
By replacing idiom A’s intermediate representation with that of idiom B, we remove any information specific to idiom A that would have been transmitted through the intermediate pathway. Figure 7(c) shows that there is a drop of $\Delta F(s_{a}\hookleftarrow s_{a}^{*})$ by $-0.09$ in layer 7. On the other hand, the probability of the literal interpretation $\Delta L(s_{a}\hookleftarrow s_{a}^{*})$ does not change significantly. This implies that the literal path is intact, and there’s a bit of noise by adding the unrelated activations which cause a slight drop of $\Delta L(s_{a}\hookleftarrow s_{a}^{*})$ compared to $\Delta L(s_{a}\hookleftarrow s_{l})$. The magnitude of the drop in $\Delta F(s_{a}\hookleftarrow s_{a}^{*})$ closely matches the decrease after patching in the literal paraphrase $\Delta F(s_{a}\hookleftarrow s_{l})$, demonstrating that the figurative interpretation is predominantly routed through the intermediate pathway. In Appendix D, we further demonstrate that ablating the attention edge from the idiom span to its subsequent token causes $\Delta L(s_{a})$ to increase while $\Delta F(s_{a})$ decreases. This provides additional evidence that the figurative meaning is predominantly conveyed through this intermediate pathway. Meanwhile, when we patch activations from idiom B into the intermediate pathway of idiom A, we expect the intermediate representations to be overwritten by those of idiom B (Figure 6(c)). This would lead to increased prediction probabilities for both the figurative and literal interpretations associated with idiom B. As shown in Figure 7(d), both $\Delta F^{B}(s_{a}\hookleftarrow s_{a}^{*})$ and $\Delta L^{B}(s_{a}\hookleftarrow s_{a}^{*})$ increase, indicating that information supporting both interpretations is transmitted through the intermediate path. However, the magnitudes of these increases are not equal, suggesting that the pathway preferentially amplifies figurative interpretation over the other.
@@ 6 Conclusion
We present multiple steps of idiom processing in a transformer model (Figure 1), and identify distinct components and flows involved in the interpretation of figurative meaning. By isolating idiomatic expressions from the context, we gain insight into which components of the model are most sensitive to figurative interpretation compared to literal one. However, contextual information plays a crucial role in interpreting idioms, as it helps determine whether a given expression should be understood in its literal or figurative sense Holsinger (2013). Future research would explore how the model’s processing changes when disambiguating context is provided. By inserting disambiguating cues both before and after the idiom, we could examine the interactions between contextual clarification and idiom comprehension.
Moreover, an important next step is to extend our analysis to a larger number of idioms and develop a diagnostic framework that predicts, from the model’s internal activations alone, whether an expression’s figurative meaning is known. Concretely, one could observe the engagement of idiom-specialized attention heads and the activation of retrieval and binding pathways for every idiom in the set. If a given idiom fails to activate its dedicated heads, this would point to a breakdown in retrieval or binding. By looking at these activation patterns, we could quantify the model’s competence with respect to the idiom comprehension. Such an approach would not only deepen our understanding of how figurative meanings are encoded in transformer representations but also provide a useful guidance for pinpointing specific processing failures in idiom comprehension.
@@
While causal tracing methods have been widely used in recent work Geva et al. (2023); Dar et al. (2022); Meng et al. (2022); Heimersheim and Nanda ; Nanda et al. (2023), they only approximate the actual information stored in activations. Moreover, knocking out or replacing the activations can lead the model to out-of-distribution behaviors and cast doubt on the robustness of any interpretability claim. Our study is limited to LLama3.2-1B-base model, which is much smaller than the current state-of-the-art transformers, scaling these analyses to larger models remains an open challenge. Also, we simplify idiom processing by excluding context, even though context critically shapes idiomatic meaning. Altogether, these factors limit the generalizability of our results.
@@ Acknowledgements
Appendix A Selected top-20 attention heads
We further extracted semantic heads $\mathcal{H}_{\mathrm{sem}}$, i.e., heads that show a positive $\Delta F(s_{a})$ and a negative $\Delta L(s_{a})$, which are processing ambiguous sentence in a literal way. Ablating them suppresses literal processing and restores figurative understanding. For comparison, we also randomly selected heads that are neither in $\mathcal{H}_{\mathrm{idiom}}$ nor $\mathcal{H}_{\mathrm{sem}}$.
Type top-$20$ attention heads (layer, head)
Idiomatic (0, 4), (0, 28), (1, 5), (2, 9), (0, 30), (0, 8), (0, 19), (2, 2), (1, 25), (1, 18), (4, 31), (1, 31), (0, 12), (0, 3), (2, 28), (2, 17), (2, 8), (2, 30), (1, 9), (4, 24)
Semantic (1, 3), (3, 28), (3, 30), (0, 10), (7, 16), (1, 20), (2, 4), (3, 12), (3, 14), (6, 27), (2, 19), (10, 11), (4, 26), (0, 18), (6, 31), (9, 22), (2, 5), (0, 9), (1, 19), (0, 11)
@@ Appendix A Selected top-20 attention heads
Table 1: Top-$20$ attention head sets for each type.
Appendix B Activation divergence analysis between idiomatic and literal contexts
To examine how idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ process idioma differently compared to literal contexts, we focus on their activations over the idiom token span. Let =$\mathcal{I}=\{i_{1},\dots,i_{K}\}\subseteq\{1,\dots,T\}$ be the indices of the idiom span tokens. Note that the literal paraphrase $s_{l}$ preserves the same idiom token span, hence $|\mathcal{I}|$ is identical for both $s_{a}$ and $s_{l}$. For each head $h\in\mathcal{H}_{\mathrm{idiom}}$, we compute the averaged activation from idiom span value vectors:
$v_{a}\;=\;\frac{1}{\mathcal{I}}\sum_{i\in\mathcal{I}}\alpha^{(a)}_{h,i}\,V_{i},\;v_{l}\;=\;\frac{1}{\mathcal{I}}\sum_{i\in\mathcal{I}}\alpha^{(l)}_{h,i}\,V_{i},$
where $\alpha^{(o)}_{h,i}$ and $\alpha^{(l)}_{h,i}$ are the attention weights on token $i$ in $s_{a}$ and $s_{l}$, respectively, and $V_{i}$ is the corresponding value projection. These mean vectors capture the direction in the head’s value subspace over the idiom span under each context. We then measure their alignment via cosine similarity:
$\cos(v_{a},\,v_{l})=\frac{v_{a}\cdot v_{l}}{\|v_{a}\|\,\|v_{l}\|}.$
A low cosine similarity between $v_{a}$ and $v_{l}$ means that the head’s representational direction shifts substantially when processing the idiomatic sentence $s_{a}$ compared to its literal paraphrase $s_{l}$, revealing its specialization in processing ambiguous sentence and extracting its figurative meaning. In contrast, a high cosine similarity indicates that the head maintains a similar representation across both contexts, reflecting a more general semantic role. For comparison, we also compute cosine similarities for the semantic heads $\mathcal{H}_{\mathrm{sem}}$ and for a random set of heads $\mathcal{H}_{\mathrm{rand}}$.
For each sentence, we compute the cosine similarity $\cos(v_{a},v_{l})$ for every head in a given set of 20 attention heads, and then average these 20 values to obtain a single mean similarity (Figure 8). The cosine similarity across the dataset is shown in Figure 8. The $cos\_random$ ($M=0.87,\,SD=0.07$), $cos\_sem$ ($M=0.80,\,SD=0.09$) is significantly larger than $cos\_idiom$ ($M=0.76,\,SD=0.12$), ($p<0.05$).
Figure 8: Cosine similarities between the attention activations on idiomatic sentences ($s_{a}$) and their literal paraphrase ($s_{l}$) for the top-20 idiomatic heads, semantic heads, and random heads.
By confirming that the heads identified as specialized in retrieving the idioms’ figurative meaning indeed have divergent attention patterns when processing idioms compared to literal paraphrases, we strengthen the evidence that these heads are responsible for the unique processing required to interpret idiomatic expressions correctly.
Appendix C Layer-wise interpretation flow through activation patching
We additionally patch idiom’s because token activation into paraphrased sentences’ because token position. The conceptual details of the experiment is described as follow.

### v4

@@ 4.2 Identifying attention heads specific to retrieving figurative interpretations
Next, we investigated which attention heads are specialized for retrieving an idiom’s figurative interpretation. We knock out attention heads iteratively at the idiom span tokens and measure $\Delta F(s_{a})$, $\Delta L(s_{a})$.
@@
Figure 3: Heatmaps of the (a) $\Delta F(s_{a})$ (b) $\Delta L(s_{a})$ when ablating individual attention heads at the idiom span. Idiomatic heads: Heads those are crucial for retrieving the figurative interpretation of idiom; $\Delta F(s_{a})$ and $\Delta L(s_{a})$.
To operationalize this, we define idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ as heads with a large negative $\Delta F(s_{a})$ and positive $\Delta L(s_{a})$. Figure 3 shows that there is a subset of attention heads that simultaneously exhibits large $\Delta F(s_{a})$ (Figure 3(a)) and $\Delta L(s_{a})$ (Figure 3(b)). To identify which of these heads are consistently important across instances of $s_{a}$, we performed a nonparametric bootstrap with $B=1000$ resamples for each head’s contribution to the interpretation. We then rank all heads by, on the one hand, $\Delta F(s_{a})$, and, on the other hand, $\Delta L(s_{a})$, and select the top 20 heads highly ranked on both orders (selected attention heads listed in Appendix Table 1).
@@ 4.3 Causal role of MLP sublayers and idiomatic heads in figurative reading
To validate that the early MLP sublayers and the selected idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ are critical for the figurative reading, we patch the activations on $s_{a}$ (at idiom span) into the corresponding components when processing $s_{l}$. As a control, we patch randomly chosen MLP sublayers and heads.
ResultsPatching these components boosts the figurative interpretation, as quantified by $\Delta F(s_{l}\hookleftarrow s_{a})$ (M=0.17,SD=0.49M=0.17,\,SD=0.49) while suppressing the literal interpretation, $\Delta L(s_{l}\hookleftarrow s_{a})$ (M=0.29,SD=0.58M=-0.29,\,SD=0.58). In contrast, patching random components barely affects the figurative interpretation ($\Delta F(s_{l}\hookleftarrow s_{a})$: M=0.008,SD=0.09M=-0.008,\,SD=0.09) as well as literal one ($\Delta L(s_{l}\hookleftarrow s_{a})$: M=0.005,SD=0.11M=-0.005,\,SD=0.11). Effects on the experiment and the control are significantly different ($p<0.05$). These contrasting effects suggest that idiomatic components are specialized to enhancing the figurative interpretation while suppressing literal interpretations, so that replacing their activations injects a figurative interpretation and amplifies $\Delta F(s_{l})$, while dropping $\Delta L(s_{l})$.
@@ 5 Tracing the interpretation flow
5.1 Locating idiom interpretation across token positions
After the retrieval process, the model encodes both literal and figurative interpretations in its residual stream. We next investigate how these two representations are weighted or integrated in the final prediction. To locate the representation of the interpretations, we calculate the mutual nearest-neighbor kernel alignment Huh et al. (2024); Cho et al. (2025) between sentence embeddings of a corresponding idiom token span from $s_{f}$ and $s_{l}$ encoded by BGE M3 Chen et al. (2024) and hidden states from various token positions of idiom sentence $s_{a}$. This metric quantifies the similarity between representations, with higher values indicating greater semantic similarity.
@@ 5.1 Locating idiom interpretation across token positions
(b) ‘because’ token
(c) After ‘because’ token
@@
ResultsWe observe that at the ‘because’ token (Figure 4(b)), the kernel alignment between the idiom’s hidden states and its figurative interpretation surpasses that of the literal interpretation after layer 4. This suggests that by this point, the model has begun to favor the figurative interpretation ( 2 selective interpretation step). Based on the results, we hypothesize an intermediate pathway where token position ‘because’ plays a pivotal role in steering the idiom toward a figurative interpretation. In parallel, to keep the alternative interpretation available at the prediction (as in Figure 4(d)), we posit a direct route that favors the literal interpretation by more strongly preserving compositional semantic information.
To measure the information flow along the intermediate pathway (through the subsequent token position ‘because’), we patch the activation $s_{a}$ at the ‘because’ token with $s_{f}$ or $s_{l}$ (Figures 5(a), 5(b)). On the other hand, patching the activation $s_{a}$ at the idiom token span with an alternative activation ($s_{f}$ or $s_{l}$) and then re-patching attention at the ‘because’ token with $s_{a}$ allows us to isolate the information that flows via the direct route (Figures 5(c), 5(d)).
[…]
@@ 5.2 Analyzing layer-wise competing interpretation flow through activation patching
Figure 6: Layer-wise interpretation shift after patching in activations. The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ exceeds the average difference across layers (paired $t$-test, $p<0.05$).
@@
ResultsFigure 6(a) shows that the patching as illustrated in Figure 5(a), increases the $\Delta F(s_{a}\hookleftarrow s_{f})$ in mid-layers 4-10 (max: $0.17$), but has a relatively small effect on $\Delta L(s_{a}\hookleftarrow s_{f})$, which shows some decrease (min: 0.09-0.09). Patching in $s_{l}$ (as in Fig. 5(b)) removes the intermediate path figurative flow, which leads to an increase in the $\Delta L(s_{a}\hookleftarrow s_{l})$ in mid-layers (6-10) as shown in Figure 6(b) (max: $0.13$). At the same time, $\Delta F(s_{a}\hookleftarrow s_{l})$ drops symmetrically in the same layers (min: 0.10-0.10). This indicates that the figurative interpretation is suppressed in favor of the literal interpretation.
We interpret this as showing that a figurative injection doesn’t lead to suppression of the literal interpretation (an asymmetry between increases and decreases). By contrast, a literal injection must first suppress the figurative interpretation in that subspace to have an effect, producing a more symmetric push–pull tradeoff. This indicates that for the mid-layers at the ‘because’ token position, the model’s residual stream is the predominant route for the figurative interpretation ( 3-1 figurative path).
Meanwhile, Figure 4(d) indicates that literal interpretation remains competitive with the figurative one. We assume that there’s a different path that favors the semantically literal interpretation ( 3-2 compositional semantic direct path) to convey its information to the final prediction, directly bypassing the intermediate path.
To probe this, we replace the idiom sentence’s idiom token span activation with the activation from a paraphrased sentence, and re-patch the ‘because’ token with the idiom sentence’s original activation (as shown in Figures 5(c), 5(d)). This preserves the intermediate pathway and isolates the contribution of the direct route. Figure 6(c) shows that removing direct path’s literal interpretation produces a large drop in $\Delta L$ (min: 0.25-0.25) and rise in $\Delta F$ (max: $0.20$). Conversely, removing the direct path’s figurative interpretation yields only a modest drop in $\Delta F$ (min: 0.10-0.10) and a small increase in $\Delta L$ (max: $0.10$). These results indicate that the direct path predominantly conveys literal information. The pattern is consistent across different models (Appendix E).
6 Localizing idiom disambiguation in context
To assess the role of context in idiom disambiguation, we’re now looking at the setup with context. We retain an instance only when the model assigns higher cumulative probability to the context-consistent candidate than to the alternative (i.e., $F(s_{a}|FC)>F(s_{a}|LC)$, $L(s_{a}|LC)>L(s_{a}|FC)$).
6.1 Probing sublayers through knockout with figurative vs. literal context
We conduct a layer-wise knockout experiment to assess how MLP and MHSA components contribute to idiom processing when the same idiomatic expressions are presented in different contexts.
(a) MLP FC
(b) MLP LC
(c) MHSA FC
(d) MHSA LC
Figure 7: Sublayer-wise interpretation shift $\Delta I$ after knockout activations at idiom span, for contexts FCFC: Figurative Context; LCLC: Literal Context.
ResultsFigure 7(a) shows that for idioms in a figurative context, knocking out the MLP in the early layers (0–3) results in a significant drop in $\Delta F$ (min: 0.62-0.62) and increase in $\Delta L$ (max: $0.66$), consistent with the pattern observed for idioms without context (Figure 2(a)). This indicates that these layers are critical for retrieving the figurative interpretation independently of context. This result is consistent across different models (Appendix F). For idioms in a literal context (7(b)), ablating the MLP layer 1 component leads to a significant drop in $\Delta F$ (min: 0.33-0.33) and an increase in $\Delta L$ (max: $0.09$), indicating that the figurative interpretation is still being retrieved. The effect persists but is small compared with the figurative context condition, where early layers at the idiom token appear to encode context that suppresses figurative retrieval while biasing toward the literal interpretation.
A similar pattern is observed in the MHSA layers. As shown in Figure 7(c), under figurative context, ablating early layers (0–3) produces a pronounced decrease in $\Delta F$ (min: 0.33-0.33) and an increase in $\Delta L$ (max: $0.48$). Under literal context, ablating the earliest layers (0–1) yields a drop in $\Delta F$ (min: 0.32-0.32), whereas only modest changes for $\Delta L$ (max: $0.16$) relative to the figurative context condition. That is, the literal context suppresses retrieval of the figurative interpretation from the very earliest layers. Moreover, ablating the layers (3–7) under literal context flips the pattern, where $\Delta F$ is increasing while $\Delta L$ is decreasing with a significant difference (Figure 7(d)). This indicates that, after the initial retrieval of figurative interpretation, the literal interpretation, which is disambiguated from context, is refined through mainly MHSA mediated integration with surrounding tokens.
In sum, the results point to a two stage context disambiguation: (a) early layers retrieve figurative evidence while beginning to encode context, followed by (b) mid-layers, MHSA-driven contextual disambiguation that resolves conflicts in favor of the context-appropriate interpretation.
6.2 Probing MHSA-mediated contextual disambiguation through KVQ patching
To examine in detail how MHSA mediates contextual disambiguation in early-mid layers, we perform KVQ patching, in which MHSA’s key–value (KV) activations are replaced while selectively preserving idiom tokens’ query (Q) activations across layers, thereby steering the model’s attention toward alternative contexts. That is, we redirect idiom token span’s attention from a figurative-context to the literal context, and vice versa. If the model’s final prediction changes from figurative to literal (or vice versa) after patching, it demonstrates that the MHSA mechanism in those specific layers is responsible for using the context to steer the disambiguation process.
[…]
ResultsAs shown in Figure 8, patching in KVs with alternative context yields a shift in interpretation of idiom. Under FC, effects appears already at layer 0, indicating immediate context sensitivity, with additional peaks at layers 2-3 and 6. Under LC, the probability of the context matched interpretation drops across layers 2–6, while the alternative interpretation rises. Taken together, these results suggest that the idiom tokens exploit contextual evidence primarily in layers 2-6, after the figurative interpretation is first retrieved in layers 0-1, with layer 0 playing a crucial early role under FC. Later layers tend to consolidate rather than redirect the interpretation.
@@ 7 Conclusion
We present a multi-stage mechanism of idiom comprehension in causal transformers, by identifying distinct components and flows for figurative and literal interpretations. Causal interventions reveal (i) early retrieval of the idiomatic interpretation, (ii) immediate use of preceding context, with later refinement when it conflicts with the retrieved interpretation, (iii) a selective pathway that carries the figurative interpretation through intermediate layers, and a bypass route that favorably delivers the literal interpretation directly to the output. To sum up, these findings link representational dynamics to contextual disambiguation, yielding a concise mechanistic interpretation of idiom processing.
@@
While causal tracing methods have been widely used in recent work Geva et al. (2023); Dar et al. (2023); Meng et al. (2022); Heimersheim and Nanda ; Nanda et al. (2023), they only approximate the actual information stored in activations. Moreover, knocking out or replacing the activations can lead the model to out-of-distribution behaviors and cast doubt on the robustness of any interpretability claim.
Moreover, we binarize idiom interpretation (figurative vs. literal), ignoring polysemy among figurative senses (e.g., ‘break the ice’ can be interpreted into initiate talk, ease tension, etc.), which can flatten nuance and bias evaluation. As a future direction, we can use soft labels for multiple figurative sense and distributional metrics.
@@ Acknowledgements
Appendix A Identified top-20 idiomatic heads
Table 1 lists the attention heads most influential in retrieving figurative meanings while suppressing literal interpretations and randomly selected heads.
Type top-2020 attention heads (layer, head)
Idiomatic (0, 4), (1, 5), (0, 30), (0, 19), (0, 8), (9, 30), (2, 2), (1, 18), (0, 21), (2, 8), (1, 9), (4, 2), (1, 24), (1, 13), (3, 18), (0, 0), (3, 24), (0, 26), (1, 27), (0, 28)
@@ Appendix A Identified top-20 idiomatic heads
Table 1: Top-2020 attention head sets.
Appendix B Data generation pipeline
We illustrate the data generation pipeline in Figure 9.
Figure 9: The pipeline of data generation.
Appendix C Layer-wise context disambiguation via logit lens
If the MHSA layers 3-7 play a causal role in context disambiguation, then we would expect that the representation at the final idiom token is fully disambiguated towards either the literal or the figurative meaning, depending on the earlier context. We employ the logit lens method nostalgebraist (2020) to measure the log-probability assigned to figurative and literal interpretations in the figurative ($FC$) vs. literal context ($LC$) conditions. At each layer, the hidden representations are projected into the output vocabulary space, yielding candidate probabilities. We then measure the difference in log-probability mass between figurative and literal candidates across layers, separately isolating contributions from the MHSA and MLP components.
(a) MLP FC-LC
(b) MHSA FC-LC
Figure 10: Layerwise differences in log-probability between figurative and literal candidates predictions under figurative (FC) vs. literal (LC) contexts, measured with the logit lens.
ResultsFigure 10 shows a clear disambiguation effect for MHSA starting around layer 8. Once MHSA starts modulating activations in a context-dependent manner, it writes this information into the residual stream, which is subsequently processed by the MLP layers. In the early to middle layers (0–9) of the MLP, the difference in log probabilities between figurative and literal candidate predictions under figurative vs. literal contexts remains minimal. However, beginning at layer 10, MLP activations begin to diverge: idioms presented in figurative contexts are increasingly encoded with figurative interpretations, whereas idioms in literal contexts are increasingly encoded with literal interpretations.
Appendix D Knockout for crucial role of early layers with other models
Results across different models in Figure 11 consistently show that early layers retrieve idioms’ figurative interpretations while suppressing literal counterparts, unlike unambiguous sentences.
Qwen2.5-0.5B
(a) MLP
(b) MHSA
(c) MLP
(d) MHSA
Llama3.1-8B
(e) MLP
(f) MHSA
(g) MLP
(h) MHSA
Qwen2.5-7B
(i) MLP
(j) MHSA
(k) MLP
(l) MHSA
Figure 11: Sublayer-wise interpretation shift $\Delta I(s)$ after ablating activations at idiom span, for sentences $s\in\{s_{a},s_{f},s_{l}\}$. Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$, $\Delta L(s_{l})$, $\Delta F(s_{f})$ with 95% confidence intervals. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant difference between $\Delta F(s_{a})$ and the others (paired $t$-test, $p<0.05$). The difference at * marked layer is larger than the average difference across all layers.
Appendix E Activation patching for information flow with other models
Across models, the intermediate and direct pathways consistently prefer different interpretations (see Figure 12).
Qwen2.5-0.5B